from 80,000 jars produced, a manufacturer samples 1,200 and finds 7 of those defective. estimate the total number of defective jars.
Answer by fractalier(6550)

Well, we use a proportion...that is we say that 7 is to 1200 as x is to 80000...
7/1200 = x/80000
now cross-multiply
1200x = 560000
and divide by 1200
x = 560000 / 1200
x = 466 2/3 or about 467 defective jars